OpenStudy (ethanjayco):

If a man earns $630 interest in one year at 7.5%, what was the amount of money invested?

OpenStudy (anonymous):

The answer is $8,400.00. To get that you would use the simple interest formula and find P, the principal amount or the investment, instead. The original formula looks like this I = Prt and after changing it your formula should look like this P = I / rt To clarify this a little I is interest, P is principal amount, R is rate, and T is time. So first thing to do is convert your rate into decimal form. To do that simply move the decimal point two numbers to the right or divide 7.5 by 100. Now put all of your known values into the equation. It should look like this now. P = 630 / ( 0.075 × 1 ) Now solve it. P = 630 / ( 0.075 × 1 ) P = 630 / 0.075 P = 8400 $8,400.00 And know you know how to use the simple interest formula to find investment! Hope this wasn't too confusing. I tried to make this as easy to understand as possible.